# Keystroke Dynamics: Concepts, Techniques, and Applications

### v2

@@ 6.3 Fusion of Keystroke Dynamics and Other Mobile Modalities
In [30], Corpus et al. described the use of accelerometer biometrics (the measurement of how a person holds his mobile device) in conjunction with keystroke dynamics. They proposed a model for security in mobile applications by combining accelerometer biometrics with previously studied keystroke dynamics. The goal was to apply their model in a real world use case. For data collection, they created an application using the IDE Android Studio 1.3.2. They used a QWERTY soft keyboard and collected the keystroke data from 30 users. Users had to type a password of 8-16 characters a total of 8 times. The collected data were stored in two different text files. The first file recorded the data related to keystroke dynamics and the second file contained the data of accelerometer biometrics. Among the 8 samples of each user, the first 6 samples were used for training and the remaining 2 samples were used for testing. Their models were - Naive Bayes, Decision Tree, J48 and Neural Networks. They developed their models on three sets of features to train their models - accelerometer biometrics only, keystroke features only and lastly a combination of the both. The neural network model showed the best performance for combined features. However, adding feature selection improved the performance. Chi Square attribute was used to rank the features and the low rank features were removed. Corpus et al. also removed the features which were highly correlated to other features. After data cleaning the accuracy increased to 68.89% after implementing the feature selection techniques. They used a pre-labeled test set to test the best model. The test accuracy was 73.33%, FAR was 27.59% and FRR was 26.67%. Corpus et al. modified their data collection tool for real world testing through a prototype. The enrollment of the username and four password samples was included in the mobile application. A web server was used for the implementation of re-modeling and testing processes. Six users participated in this test. These was no password sample of these six users in the database. Each user chose their own password and typed the password four times. The users tested their own passwords and the passwords of their co-participants. For the prototype combining accelerometer biometrics with keystroke dynamics, they concluded that accelerometer biometrics has an impact on keystroke dynamics based authentication. Though the prototype showed low FAR (7%), the recognition performance was above average (60%-70%). It concluded that the prototype was not fully accurate and ready to deploy.
@@
In [70], Lee et al. tried to analyze smartphone based keystroke dynamics using a 6-digit pin. Lee et al. used distance-based classification algorithm and extracted two types of smartphone data - motion and keystroke data. The extracted keystroke features included time of events, size of the fingertip, coordinates of the touch point etc. The extracted motion features included accelerometer, rotation, gravity etc. The raw data was collected by typing the 6-digit PIN - "766420". The app was installed on a Nexus 5X. Data was collected from 22 users with 100 samples from each user. One sample contained one time input of 6-digit PIN. They used two algorithms for their experiments - distance-based classification and OCSVM (one-class SVM). In the case of distance-based classification, two scaling methods were used - MinMax scaling and standard scaling. Two distance metrics were used - Euclidean distance and Manhattan distance. The distance-based classification had two results.
@@
    Adding features from motion data: Lee et al. used five different formulas featuring motion data for their experiment. The best result was achieved using standard scaling and Manhattan distance. The lowest EER was 12.63%. The best result was achieved using only "mean" formula. Lee et al. obtained an EER of 8.94% using only the keystroke data. The EER was decreased by 1.05% to 7.89% when the motion information was introduced.
@@
    FAR by Gender Match: It was observed from the result that opposite gender’s match had an impact on FAR. Lower FAR was obtained if the gender of a genuine user and the gender of an impostor were different. For male genuine users, the FAR decreased by 4.07% in case of different gender. For female genuine users, the FAR decreased by 0.64% in the case of male impostors.
@@ 7 Other Applications of Keystroke Dynamics
The number of applications that rely on input from a computer keyboard is increasing. Computers have become more commonplace in the home, workplace, and academia. People have become more comfortable with typing on keyboards. In fact, today’s self-taught typists type almost as fast as touch typists [75]. This increase in the number of people that are typing on a keyboard creates additional opportunities to apply keystroke dynamics to domains not previously envisioned. An authentication API for adopting keystroke authentication is typingdna. This keystroke authentication API can be used to secure logins, enforce reset passwords, and authenticate subjects.
@@ 7.1 Educational Level Classification
Tsimperidis et al. [121] used keystroke dynamics for the detection of one’s level of education. They proposed the randomized radial basis function network which was a novel machine learning model. Their model was capable of identifying the educational level of the user of the keyboard. Their proposed model was the first model leveraging keystroke biometrics for detection of ones education level. Their experiment was performed in two steps. First, free text unrestricted data was collected from the participants for ten months. The second step was developing the R2BN model. The computational time of the R2BN model was much higher. In order to reduce that time, they decreased the size of the dataset without degrading the performance of the model. A filter-based random subfield data condensation approach was proposed. 242 log files were collected from five educational level classes. There were 2800 to 4500 keystrokes in each log file. 162 features were extracted from the data. The data was fed into the R2BN model. This model was developed by the randomized modification of a popular radial basis function neural network (RBFN) proposed by Broomhead and Lowe [21]. The advantage of R2BN model was its speedy convergence and smaller extrapolation errors due to a shallow-learning architecture. This model achieved more than 85% accuracy in the educational level classification task using keystroke dynamics.
@@ 7.2 Emotion Detection
Detecting emotions from your counterpart during electronic communications such as email, or chat is often difficult. Written words can be misinterpreted into thinking the other side is angry, or conversely saddened, when in reality, the real emotion behind the discussion is unknown. Determining the emotion associated with the text can lead to improved communications. Feelings such as anger, sadness, disgust, fatigue, joy, fear and indifferent can be assumed from the way a user types [86]. Associating emotions based on the way an individual types can reduce guesswork on the emotional state of the author, and reduce confusion. Such a system can also serve as a warning for users to rethink or delay impulsively sending emails or other messages when angry, preventing regret.
@@
Qi, Jia, and Gao [93] proposed an emotion recognition system based on piezoelectric keystroke dynamics. Piezoelectric materials are able to reflect the force detection sensitivity. Their experiment was simply password entry. They used time and pressure dimension features. For emotion recognition, they utilized the discrete PAD 3dimensional-model. International Affective Digitized Sounds (IADS) (a standardized set of 110 digitalized sounds) was implemented to reveal emotions. In order to measure the amount of emotion induction, they implemented a PAD emotion scale. Their target was to classify 4 emotions - happiness, fear, sadness, disgust. They obtained an accuracy of 78.31% using the Random Forest Classifier. Their proposed approach proved to be a pragmatic one for emotion classification in practical applications.
@@
Nahin et al. [86] tried to identify the emotions of users by analyzing textual patterns of the user’s keystroke dynamics. They considered seven basic emotional classes - anger, guilt, disgust, joy, fear, shame and sadness. They used the categories from the International Survey on Emotion Antecedents and Reactions (ISEAR) [102]. There were two steps in their data collection process. 1) They used Java-based software to collect fixed-text data, 2) An app written in C# was used to collect free-text data. After data collection, a software program written in C was used for keystroke feature extraction from the log files in order to separate feature files. 19 keystroke features were extracted from the collected data. Nahin et al. used the best seven features for analysis. Weka 3.6.6 was used for training. Nahin et al. developed different models for fixed-text and free-text data. Testing was done using these models. For text pattern analysis, ISEAR dataset was used [101]. In order to determine emotion class from raw text, they used VSM or term vector model [99]. For fixed-text analysis, Weka tool was used [127]. Their collected data were used in order to train and develop models. These models were used to evaluate test datasets. They used these algorithms - simple logistics, SMO, Multilayer Perceptron, Random Tree, J48 and BF Tree. Nahin et al. used these same methods for free-text analysis. They calculated the accuracies of emotion class recognition for both fixed-text and free-text. In the case of VSM, the raw data were used as input. VSM identified the emotional state of a sentence by analyzing the text. In the case of the final result, only those results were considered right which were same for both keystroke dynamics and text pattern analysis. The combined results showed better accuracies than separate results.
7.3 Age, Gender and Demographics Prediction

### v3

@@ 6 Keystroke Authentication on Touch Screen and Mobile Devices
6.1 Impact of Device Models and Orientations
El-Abed, Dafer, and Rosenberger [38] proposed a benchmark for keystroke dynamics using a mobile phone and tablet. The touch screen-based benchmark was developed in both portrait and landscape orientations. The objective of this benchmark was to evaluate the keystroke authentication algorithms for varying models of devices and different orientations. An online visualizer was provided to observe the acquired keystroke signals. They used an Android application for developing their benchmark. They used the Nexus 5 and the Samsung Galaxy Note 10.1 2014 tablet for data collection. There were two main purposes of this keystroke benchmark. Firstly, they wanted to investigate the patterns of typing of the users for a variety of orientations, but the device and the password were the same. Secondly, they wanted to analyze the effect of different devices on the typing patterns of the users, keeping the orientation and the password constant. Data was gathered from 47 users. The users type the password “rhu.university” in four configurations in each session: phone/portrait, tablet/portrait, phone/landscape, tablet/landscape. The key features of the benchmark were: PP, PR, RP, RR, TT (time of typing the password), Screen Orintation (portrait or landscape), and Screen Size. They then performed a statistical analysis using their benchmark and used the Kruskal-Wallis test [55] for the analysis. The goal was to analyze the changes in the keystroke features due to different orientations and different devices. The results indicated that the keystroke features varied significantly for different orientations and different devices. These results must be considered for touch-screen-based mobile authentication technology.
6.2 Impact of Different Types of Features
6.2.1 Combination of Touch-Based and Timing Features
Gautam and Dawadi [42] analyzed different features of the touch-screen keyboard for keystroke authentication. They combined the effect of touch-based and timing features and developed a dataset of 7 users. The authentication system consisted of three steps: Enroll, Verify, and Identify. The user authorization process consisted of two phases. For both the training and enrollment phases, a reference template was generated. In the testing or authentication phase, the test samples were matched against the reference template using the Median Vector Classifier [6]. They used four features for authentication: Key hold time (H), Flight time (FT), Pressure (P), Area (A). They used the string .pie7Crawl for their dataset. The dataset consisted of 47 features. In order to perform their experiment, Gautam and Dawadi utilized a OnePlus 3 Android phone with an AMOLED capacitive touch screen and 1080 * 1920 pixels (5.5 inches). All 7 participants were touch screen smartphone users. Every user typed the string .pie7Crawl7 on the touch screen keyboard. Among the 7 participants, one was considered a legal user, and the other 6 participants were considered imposters. The legal user participated in 40 sessions. Among those sessions, 10 were reserved for training and the rest for testing. Each of the other 6 participants (imposters) participated in 5 sessions. By using the median vector proximity algorithm, they achieved an average EER of 8.33% and a standard deviation EER of 7.07%.
6.2.2 Interaction of Touch Screen Features
Antal, Szabo, and Laszlo [15] used their dataset of 42 individuals to investigate the interaction of touchscreen features on the performance of keystroke authentication systems. For data collection, they developed an Android application for a Nexus 7 tablet and a Mobil LG Optimus L7 P710. The application had its own software keyboard. During the registration period, the users had to enter their data, such as their gender, date of birth, experience of smartphone usage, etc. Data was collected in multiple sessions. Most users participated in two sessions in two weeks. The users entered the same password (tie5Roanl) during each session 30 times. A total of 42 individuals provided their data, with 51 entries for each user. Entries containing deletions were discarded from the dataset. As all users entered the same password, the data of each user could be used both as a legal user and as an impostor. Among the 42 users, 37 were tablet users and 5 were mobile phone users. The system saved both the timing and touchscreen features. They used WEKA (version 3.6.11) [88] software for their experiment. Weka’s search methods were used to optimize some of the default parameters of the classifiers. They used Bayesian network, Nearest neighbors (k-NN, IBk in Weka), Naive Bayes, Decision trees, Support vector machines, Multilayer perceptrons, etc., for the evaluation process.
In the identification phase, Antal, Szabo, and Laszlo used two datasets to compute the accuracy, one with touch screen features (pressure and finger area) and one without touchscreen features. There were 41 features in the first one and 71 features in the second. The data was used without any modification or feature selection and without boosting or tuning methods. They used 10 fold cross-validation in their experiment and computed the classification accuracies by taking the average of 10 runs. All algorithms performed better on the dataset containing emphtouch screen features. In the verification phase, an R script developed by Killourhy and Maxion [66] was used for performing the measurements. Three anomaly-detection algorithms were included in the R script: Euclidean, Manhattan, and Mahalanobis metrics. The normalized data was split into three equal segments. In each part, there were 17 samples from each user. The user’s template was generated using two-thirds of the data, with the remaining data used for FRR testing. In the case of FAR testing, the first five samples from each user were used as impostor data. The measurements were performed three times, using a different fold each time for training and testing. The lowest EER was 12.9% which was achieved using the Manhattan metric for both timing and touchscreen features. For timing features, the lowest EER was 15.3%, achieved by the Manhattan metric. Their findings showed that in the identification phase, the accuracy of each classifier increased by 10% due to the addition of touch screen features. In the verification phase, the EER was reduced by 2.4%.
6.2.3 Two New Features - Speed and Distance
Kambourakis et al. [63] developed and evaluated a touchstroke-based authentication system in the Android platform. They used two new features for their experiment apart from common keystroke features viz, speed and distance. They adopted two commonly used authentication procedures. In the first procedure, the users had to type a password, and in the second case, the users had to type a passphrase. 20 users participated in the data collection between the ages of 19 and 21 years. All users had a touch screen based Android smartphone. Kambourakis et al. implemented a prototype of their proposed touchstroke-based authentication system in Google’s Android OS. In order to generate the typing behavioral profile of the user and authenticate the user, they used Waikato Environment for Knowledge Analysis (Weka) as the classification engine. There were two main analysis sub-systems in their mechanism. 1) Enrollment: Here, the typing profile of a user was created and used to train the classifiers. 2) Authentication: Here, the legitimacy of a user was checked. The typing profile of a user was saved in a database. They considered three machine learning algorithms for their experiment: Random Forest, k-NN, and MLP. However, MLP was rejected due to the memory limitation of smartphones.
For the password scenario, Kambourakis et al. achieved the best results using the Random Forest classifier and the first data analysis procedure. For the passphrase scenario, k-NN performed better using the second data analysis procedure. For the password scenario, the average FAR%, FRR%, and EERvalues were 12.5, 39.4, and 26, respectively. For the passphrase scenario, the average FAR%, FRR%, and EERwere 23.7, 3.5, and 13.6, respectively. They performed their experiment on a Sony Ericsson Xperia ray. The device had a 1 GHz CPU, 512 MB RAM, a 3.3-inch touchscreen, and the Ice Cream Sandwich Android OS. This percentage varied between 91% and 100% at the time of training. The results showed that the first data analysis procedure needed more memory to perform, as it generated more data than the second one. Both algorithms were able to perform directly on the device, and it took less than one second to authenticate the user.
6.2.4 Novel Sensor Features
Giuffrida et al. [48] developed a novel keystroke authentication system enhanced with sensor data for authentication of mobile devices. They used novel sensor features to analyze the typing behavior of a user. Machine learning techniques were used to authenticate users and to associate the timing-based keystroke features with movement sensor information. They developed UNAGI, a fixed-text authentication system for Android. Their current prototype modified the traditional Android software keyboard and leveraged a number of modules, e.g., feature extraction, training, and detection. The sensor-enhanced keystroke dynamics techniques were implemented through these support modules. The user had to type a fixed-text password during an authentication session. Their authentication system processed this password for analysis. When the user typed, UNAGI recorded the key-press events and also sampled movement sensor data periodically. They used two Android sensor sampling interfaces: TYPE_LINEAR_ACCELERATION and TYPE_GYROSCOPE. Sensor values were collected at a high sampling frequency. In order to accomplish this, the SENSOR_DELAY_FASTEST flag was specified at sensor listener registration time. The feature extraction module of UNAGI processed the collected data and extracted suitable features. Features from the training sessions were fed to the training module to develop a user profile. The detection module compared the features of a testing session against the user profiles for authentication of genuine users. The movement sensor features considered by UNAGI were root mean square, number of local maxima and minima, minimal and maximal value, mean delta, the sum of negative values, the sum of positive values, mean value, mean value during KD and KU events, and standard deviation. The keystroke features considered by UNAGI were KD-KD time, KD-KU time, KU-KD time, and KU-KU time. After the feature extraction phase, the output was a vector that consisted of all the features, both keystroke and movement sensor features. In their experiment, threshold-based binary classification algorithms were suitable. Giuffrida et al. used one-class SVM, Naive Bayes, k-Nearest Neighbours (kNN), and the mean algorithm. They also used several distance metrics - Euclidean, Euclidean normed, Manhattan, Manhattan scaled, and Mahalanobis. They used three different configurations for evaluation purposes: keystroke data only, sensor data only, and a combination of both. Results showed better accuracy with sensor-based features rather than keystroke timing features, i.e., an EER of 0.08% for sensor-based features vs. an EER of 4.97% for keystroke timing features.
6.2.5 Touch-Based Features with Three Distinct Hand Postures
Buschek, De Luca, and Alt [22] conducted an experiment investigating mobile keystroke biometrics. They compared touch-based features in consideration of three distinct hand postures and evaluated touch-specific features for user authentication. They used two types of typing features:
  1. 1.
    Temporal features: hold time, flight time, up-up times, down-down times, etc.
  2. 2.
    Spatial touch-specific features: exact touch locations, offsets, touch jumps, drag, touch area sizes, ellipse axes, touch pressure, etc.
They used two types of models: 1) anomaly detectors, and 2) classification methods. For this study, data were collected in two sessions with a one-week gap between sessions. A total 28 users participated in the data collection, and the average age of the users was 25.
A Nexus 5 phone was utilized for the data collection. A unique software-based keyboard was used to extract the features. Each user participated in two sessions. Each session included data collected using three separate hand orientations. The participants were instructed to type 6 distinct passwords 20 times using each hand posture. From their analysis, it was observed that models achieved better results on data from a single session - the EERs were less than half of the values obtained for data across sessions. In the case of evaluation, classifiers performed better than anomaly detectors (11.7-31.1% lower EERs). EERs increased by 86.3% for different hand postures. Regarding the features, the best spatial feature sets performed better than the best temporal sets. EERs were reduced by 14.3 - 23.5% using the spatial touch features relative to the traditional temporal features. Also, EERs were reduced by using spatial touch features instead of pressure features. The combination of spatial and temporal features provided the best results; EERs were reduced by 8.5 - 26.3% relative to the best spatial features and by 26.4 - 36.8% relative to the best temporal features.
@@ 6.3 Fusion of Keystroke Dynamics and Other Mobile Modalities
Corpus et al. [30] described the use of accelerometer biometrics (the measurement of how a person holds their mobile device) in conjunction with keystroke dynamics. They proposed a model for security in mobile applications by combining accelerometer biometrics with previously studied keystroke dynamics. The goal was to apply their model in a real-world use case. For data collection, they created an application using Android Studio 1.3.2. They used a QWERTY soft keyboard and collected the keystroke data from 30 users. Users had to type a password of 8-16 characters a total of 8 times. The collected data were stored in two different text files. The first file recorded the data related to keystroke dynamics, and the second file contained the data on accelerometer biometrics.
Among the 8 samples of each user, the first 6 samples were used for training and the remaining 2 samples were used for testing. Their models were: Naive Bayes, Decision Tree, J48, and Neural Networks. They developed their models on three sets of features to train their models - accelerometer biometrics only, keystroke features only, and lastly, a combination of both. The neural network model showed the best performance for combined features. However, adding feature selection improved the performance. Chi Square attribute was used to rank the features, and the low-rank features were removed. Corpus et al. also removed features that were highly correlated with other features. After data cleaning, the accuracy increased to 68.89% by implementing the feature selection techniques. They used a pre-labeled test set to test the best model. The test accuracy was 73.33%, FAR was 27.59% and FRR was 26.67%. Corpus et al. modified their data collection tool for real-world testing through a prototype. The enrollment of the username and four password samples was included in the mobile application. A web server was used to implement re-modeling and testing processes. Six users participated in this test. There was no password sample for these six users in the database. Each user chose their own password and typed the password four times. The users tested their own passwords and the passwords of their co-participants. For the prototype combining accelerometer biometrics with keystroke dynamics, they concluded that accelerometer biometrics has an impact on keystroke dynamics-based authentication. Though the prototype showed low FAR (7%), the recognition performance was above average (60%-70%). It concluded that the prototype was not fully accurate and ready to deploy.
@@
Lee et al. [70] analyzed smartphone-based keystroke dynamics using a 6-digit pin. They used a distance-based classification algorithm and extracted two types of smartphone data, motion and keystroke data. The extracted keystroke features included the time of events, size of the fingertip, coordinates of the touch point, etc. The extracted motion features included accelerometer, rotation, gravity, etc. The raw data was collected by typing the 6-digit PIN - “766420”. The app was installed on a Nexus 5X. Data was collected from 22 users with 100 samples from each user. One sample contained a one-time input of a 6-digit PIN. They used two algorithms for their experiments, distance-based classification and OCSVM (one-class SVM). In the case of distance-based classification, two scaling methods were used: MinMax scaling and standard scaling. Two distance metrics were used: Euclidean distance and Manhattan distance. The distance-based classification experiments had two results:
@@
    Adding features from motion data: Lee et al. used five different formulas featuring motion data for their experiment. The best result was achieved using standard scaling and Manhattan distance. The lowest EER was 12.63%. Lee et al. obtained an EER of 8.94% using only the keystroke data. The EER was decreased by 1.05% to 7.89% when the motion information was introduced.
@@
    FAR by Gender Match: It was observed from the result that the opposite gender’s match had an impact on FAR. A lower FAR was obtained if the gender of a genuine user and the gender of an impostor were different. For male genuine users, the FAR decreased by 4.07% in the case of different genders. For female genuine users, the FAR decreased by 0.64% in the case of male impostors.
@@ 7 Other Applications of Keystroke Dynamics
The number of applications that rely on input from a computer keyboard is increasing as computers have become more commonplace in the home, workplace, and academia. People have become more comfortable with typing on keyboards. In fact, today’s self-taught typists type almost as fast as touch typists [75]. This increase in the number of people who are typing on a keyboard creates additional opportunities to apply keystroke dynamics to domains not previously envisioned. An authentication API for adopting keystroke authentication is typingdna. This keystroke authentication API can be used to secure logins, enforce password resets, and authenticate subjects.
@@ 7.1 Educational Level Classification
Tsimperidis et al. [121] used keystroke dynamics for the detection of one’s level of education. They proposed the randomized radial basis function network, which is a novel machine learning model. Their model is capable of identifying the educational level of the keyboard user. Their model is the first model leveraging keystroke biometrics for the detection of one’s education level. Their experiment was performed in two steps. First, free-text unrestricted data were collected from the participants for ten months. The second step was developing the R2BN model. The computational time of the R2BN model was very high. In order to reduce that time, they decreased the size of the dataset without degrading the performance of the model. A filter-based random subfield data condensation approach was proposed. 242 log files were collected from five educational-level classes. There were 2800 to 4500 keystrokes in each log file. 162 features were extracted from the data. The data was fed into the R2BN model. This model was developed by the randomized modification of a popular radial basis function neural network (RBFN), proposed by Broomhead and Lowe [21]. The advantage of R2BN model was its speedy convergence and smaller extrapolation errors due to a shallow-learning architecture. This model achieved more than 85% accuracy in the educational level classification task using keystroke dynamics.
@@ 7.2 Emotion Detection
Detecting emotions from your counterpart during electronic communications, such as email or chat, is often difficult. Written words can be misinterpreted into thinking the other side is angry, or conversely, saddened, when in reality, the real emotion behind the discussion is unknown. Determining the emotion associated with the text can lead to improved communication. Feelings such as anger, sadness, disgust, fatigue, joy, fear, and indifference can be assumed from the way a user types [86]. Associating emotions based on the way an individual types can reduce guesswork on the emotional state of the author and reduce confusion. Such a system can also serve as a warning for users to rethink or delay impulsively sending emails or other messages when angry, preventing regret.
@@
Qi, Jia, and Gao [93] proposed an emotion recognition system based on piezoelectric keystroke dynamics. Piezoelectric materials are able to reflect the force detection sensitivity. Their experiment was simply a password entry. They used time and pressure dimension features. For emotion recognition, they utilized the discrete PAD 3-dimensional model. International Affective Digitized Sounds (IADS) (a standardized set of 110 digitalized sounds) was implemented to reveal emotions. In order to measure the amount of emotion induction, they implemented a PAD emotion scale. Their target was to classify 4 emotions - happiness, fear, sadness, and disgust. They obtained an accuracy of 78.31% using the Random Forest Classifier. Their proposed approach proved to be a pragmatic one for emotion classification in practical applications.
@@
Nahin et al. [86] identified the emotions of users by analyzing the textual patterns of the user’s keystroke dynamics. They considered seven basic emotional classes, viz. anger, guilt, disgust, joy, fear, shame, and sadness. They used the categories from the International Survey on Emotion Antecedents and Reactions (ISEAR) [102]. There were two steps in their data collection process. 1) A Java-based software to collect fixed-text data, 2) An app written in C# to collect free-text data. After data collection, a software program written in C was used for keystroke feature extraction from the log files in order to separate feature files. 19 keystroke features were extracted from the collected data. Nahin et al. used the best seven features for analysis. Weka 3.6.6 was used for training. They developed different models for fixed-text and free-text data. For text pattern analysis, the ISEAR dataset was used [101]. In order to determine emotion class from raw text, they used VSM or term vector model [99]. For fixed-text analysis, the Weka tool was used [127]. These models were used to evaluate test datasets. They used these simple logistics, SMO, Multilayer Perceptron, Random Tree, J48, and BF Tree for evaluation. They used these same methods for free-text analysis. They calculated the accuracies of emotion class recognition for both fixed-text and free-text. In the case of VSM, the raw data were used as input. VSM identified the emotional state of a sentence by analyzing the text. In the case of the final result, only those results were considered right that were the same for both keystroke dynamics and text pattern analysis. The combined results showed better accuracies than the separate results.
7.3 Age, Gender, and Demographics Prediction